GoodFellow:

Consecutive standing waves occur at frequencies 280 Hz and 315 Hz on a taut string fixed at both ends. Determine the frequency of the 4th harmonic.

Vocaloid:

Nth harmonic frequency formula for a string w/ fixed ends: fn = (nv)/(2L) We have two consecutive frequencies but we don’t know what n-value either of them are. So we let the lower frequency 280 hz correspond to n and the higher frequency correspond to n + 1 giving us 280 = nv(2L) 315 = (n+1)v/(2L) Divide the two equations to eliminate L and v. Solve for n. Once you have your n value, you know that n * (f1) = 280 Solve for the fundamental frequency f1. Finally, multiply 4 * f1 to find the fourth harmonic.

Vocaloid:

correction: 280 = nv/(2L) with division between the nv and the 2L

GoodFellow:

Good Job Vocaloid. I knew the answer. For your understanding, I would say, (this is my way) just expand the second equation. Then you get the term corresponding to 280 in it. Just go ahead and subtract & get the value of\[\frac{v}{2L}\]. To get the final answer just multiply that by 4. Do not take effort to find n. Good Job! again.